### v1

@@ Abstract
Background and Objectives: Monitoring sleep states is crucial in modern medical research for assessing sleep quality and diagnosing sleep disorders. Traditional manual staging methods are not only time-consuming but also subject to subjective judgment, leading to inconsistent results. This study developed an automated sleep staging and sleep disorder classification model through deep learning technology, aimed at improving diagnostic accuracy and efficiency.
Methods: Considering the characteristics of polysomnography (PSG) multi-lead sleep monitoring, we designed a multimodal sleep state classification model, MSSC-BiMamba, that combines an Efficient Channel Attention (ECA) mechanism with a Bidirectional State Space Model (BSSM). The ECA module allows for weighting data from different sensor channels, thereby amplifying the influence of diverse sensor inputs. Additionally, the implementation of the bidirectional mamba structure enables the model to effectively capture the multidimensional features and long-range dependencies of PSG data.
Results: The developed model demonstrated impressive performance on sleep stage classification tasks. On the ISRUC-S3 dataset, the model achieved an accuracy of 0.852, an F1 score of 0.824, and a Cohen’s kappa score of 0.803. On the ISRUC-S1 dataset, the model attained scores of 0.830 for accuracy, 0.801 for F1 score, and 0.773 for Cohen’s kappa. Furthermore, the model exhibited an accuracy of 0.952 for sleep health prediction when evaluated on a combined dataset consisting of ISRUC and Sleep-EDF.
Conclusion: Our model is the first to apply the bidirectional Mamba to sleep staging with complex PSG data, showing substantial gains in computational and memory efficiency over traditional Transformer-style models. This method not only makes health monitoring more accessible but also broadens the reach of advanced healthcare, thereby enhancing sleep health management with innovative technology.
@@ 1 Introduction
Sleep, a fundamental aspect of human health, involves multiple stages that play critical roles in cognitive function and overall well-being [1]. The accurate classification of sleep stages and assessment of sleep health are vital for diagnosing and treating sleep disorders. Traditionally, polysomnography (PSG) has been the gold standard for monitoring sleep stages; however, it requires specialized equipment and is generally confined to clinic settings. This limitation underscores the need for more accessible and less intrusive methods of sleep analysis.
@@
  • Innovative Model Architecture: We introduce the MSSC-BiMamba model, which combines the Efficient Channel Attention mechanism with the Bi_Mamba deep learning architecture, tailored specifically for sleep stage classification and health status discrimination.
  • Enhanced Efficiency: The MSSC-BiMamba model, employing multimodal PSG data, significantly improves computational and memory efficiency, bridges the gap between intricate clinical evaluations and practical, real-time monitoring solutions.
@@ 4.1 Classification experiments on the ISRUC-S3 and ISRUC-S1 datasets
Therefore, our model can effectively automate sleep stage classification based on clinical PSG data, freeing clinical physicians from the tedious task of manual staging. This allows them to focus more on supervising and correcting the automated classifications, as well as devoting more energy to the treatment and care of patients. This automated classification method not only improves diagnostic efficiency but also enhances accuracy when dealing with large volumes of data, bringing significant technological innovations to the field of sleep medicine.

### v2

@@ Abstract
Monitoring sleep states is essential for evaluating sleep quality and diagnosing sleep disorders. Traditional manual staging is time-consuming and prone to subjective bias, often resulting in inconsistent outcomes. Here, we developed an automated model for sleep staging and disorder classification to enhance diagnostic accuracy and efficiency. Considering the characteristics of polysomnography (PSG) multi-lead sleep monitoring, we designed a multimodal sleep state classification model, MSSC-BiMamba, that combines an Efficient Channel Attention (ECA) mechanism with a Bidirectional State Space Model (BSSM). The ECA module allows for weighting data from different sensor channels, thereby amplifying the influence of diverse sensor inputs. Additionally, the implementation of bidirectional Mamba (BiMamba) enables the model to effectively capture the multidimensional features and long-range dependencies of PSG data. The developed model demonstrated impressive performance on sleep stage classification tasks on both the ISRUC-S3 and ISRUC-S1 datasets, respectively containing data with healthy and unhealthy sleep patterns. Also, the model exhibited a high accuracy for sleep health prediction when evaluated on a combined dataset consisting of ISRUC and Sleep-EDF. Our model, which can effectively handle diverse sleep conditions, is the first to apply BiMamba to sleep staging with multimodal PSG data, showing substantial gains in computational and memory efficiency over traditional Transformer-style models. This method enhances sleep health management by making monitoring more accessible and extending advanced healthcare through innovative technology.
@@ 1 Introduction
Sleep, a fundamental aspect of human health, involves multiple critical stages in cognitive function and overall well-being [1]. The accurate classification of sleep stages and assessment of sleep health are vital for diagnosing and treating sleep disorders. Traditionally, polysomnography (PSG) has been the gold standard for monitoring sleep stages; however, it requires specialized equipment and is generally confined to clinic settings. This limitation underscores the need for more accessible and less intrusive methods of sleep analysis.
@@
  • Innovative Model Architecture: We introduce the MSSC-BiMamba model with an innovative architecture, which combines the Efficient Channel Attention mechanism with BiMamba, tailored specifically for sleep stage classification and health status discrimination.
  • Enhanced Efficiency: The MSSC-BiMamba model, employing complex multimodal PSG data, significantly improves computational and memory efficiency, and bridges the gap between intricate clinical evaluations and practical, real-time monitoring solutions.
@@ 4.1 Classification experiments on the ISRUC-S3 and ISRUC-S1 datasets
Therefore, the developed model demonstrated impressive performance on sleep stage classification tasks on both the ISRUC-S3 and ISRUC-S1 datasets. These datasets respectively contain clinical PSG data from populations with healthy and unhealthy sleep patterns. This distinction highlights the model’s capability to effectively differentiate and analyze sleep stages across diverse health conditions, showcasing its potential for broad applications in sleep medicine. Especially, our efficient model can free clinical physicians from the tedious task of manual staging, allowing them to focus more on supervising and correcting the automated classifications, as well as devoting more energy to the treatment and care of patients. This automated classification method not only enhances diagnostic accuracy, but also significantly improves efficiency when dealing with large volumes of data, bringing significant technological innovations to the field of sleep medicine.